# A Computational Theory for Efficient Model Evaluation with Causal Guarantees

### v3

@@ Appendix D Hyper-parameters of evaluation model
Appendix E Original evaluation errors
We listed the original average root mean square evaluation errors for 4 different evaluation metrics: ROC-AUC, ACC, RMSE, $R^{2}$ in the 11 scenes for evaluation model learning.
@@ Appendix E Original evaluation errors
Linear 0.00474 0.00918 0.00027 0.00077 0.00555
MLP 0.03528 0.31249 0.02601 0.05205 0.03862
SVR 0.03881 0.04573 0.03729 0.02167 0.03665
RF 0.01348 0.00513 0.00193 0.00335 0.01378
LGBM 0.00998 0.00530 0.00219 0.00301 0.00985
XGBoost 0.01223 0.00549 0.00235 0.00331 0.01183
CatBoost 0.00756 0.00468 0.00159 0.00241 0.00695
@@
Linear 0.01689 0.00061 0.00061 0.00174 0.01208
MLP 0.04252 0.02690 0.02690 0.05219 0.04072
SVR 0.06077 0.04628 0.04628 0.03578 0.05020
RF 0.04748 0.00441 0.00441 0.00770 0.03020
LGBM 0.03539 0.00508 0.00508 0.00702 0.02145
XGBoost 0.04258 0.00521 0.00521 0.00756 0.02537
CatBoost 0.02669 0.00368 0.00368 0.00553 0.01515
@@
Linear 0.01587 0.01438 0.00723 0.01148 0.05123 0.01391
MLP 0.04397 0.04000 0.03974 0.05053 0.06654 0.04581
SVR 0.03433 0.02250 0.05526 0.03135 0.06141 0.07091
RF 0.02107 0.01608 0.01014 0.01645 0.05892 0.02255
LGBM 0.01934 0.01517 0.01056 0.01407 0.05555 0.01890
XGBoost 0.02116 0.01635 0.01016 0.01595 0.05970 0.02229
CatBoost 0.01748 0.01432 0.00888 0.01262 0.05232 0.01839
@@
Linear 0.01850 0.01438 0.00696 0.01887 0.03357 0.01391
MLP 0.04886 0.03994 0.03924 0.05166 0.06719 0.04581
SVR 0.07261 0.03029 0.03749 0.07211 0.06634 0.07091
RF 0.03444 0.01641 0.00973 0.03068 0.05216 0.02255
LGBM 0.02929 0.01526 0.00972 0.02610 0.04359 0.02255
XGBoost 0.03395 0.01656 0.01000 0.02981 0.05219 0.01890
CatBoost 0.02655 0.01443 0.00879 0.02426 0.04645 0.01839
Appendix F Evaluation cost
We list the table to compare the average evaluation time of experimental (simulation-based) evaluation and our computation evaluation as following table 12.
@@ Appendix F Evaluation cost
Table 12: Expected evaluation time comparison in 12 scenes per subject.“S” means simulation.
Appendix G Upper bounds tables
@@
Appendix H Paradigm of computational evaluation
@@
Figure 6: The decision tree that how to use computational evaluation models in your own evaluation task. The GEE means generalized evaluation error and the GCEE means generalized causal evaluation error. EC is evaluation condition, S is evaluation subject, and M is evaluation metric. P(M(S)|EC) is the potential outcome of evaluation metric after deploying the subject given evaluation condition.

### v4

@@ Appendix F Hyper-parameters of evaluation model
Appendix G Original evaluation errors
We listed the original average root mean square evaluation errors for 4 different evaluation metrics: ROC-AUC, ACC, RMSE, $R^{2}$ in the 11 scenes for evaluation model learning. The reason we do not use the policy risk, Area Under the Qini Curve (Qini Coefficient, Radcliffe [2007]) and Area Under the Uplift Curve (AUCC, Radcliffe [2007]) for discrete target is that not all models return a probability for classification. For example, the output continuous value of logistic regression can not be used to represent the probability that it belong to a class.
@@ Appendix G Original evaluation errors
Het-Linear 0.00474 0.00918 0.00027 0.00077 0.00555
Het-MLP 0.03528 0.31249 0.02601 0.05205 0.03862
Het-SVR 0.03881 0.04573 0.03729 0.02167 0.03665
Het-RF 0.01348 0.00513 0.00193 0.00335 0.01378
Het-LGBM 0.00998 0.00530 0.00219 0.00301 0.00985
Het-XGBoost 0.01223 0.00549 0.00235 0.00331 0.01183
Het-CatBoost 0.00756 0.00468 0.00159 0.00241 0.00695
@@
Het-Linear 0.01689 0.00061 0.00061 0.00174 0.01208
Het-MLP 0.04252 0.02690 0.02690 0.05219 0.04072
Het-SVR 0.06077 0.04628 0.04628 0.03578 0.05020
Het-RF 0.04748 0.00441 0.00441 0.00770 0.03020
Het-LGBM 0.03539 0.00508 0.00508 0.00702 0.02145
Het-XGBoost 0.04258 0.00521 0.00521 0.00756 0.02537
Het-CatBoost 0.02669 0.00368 0.00368 0.00553 0.01515
@@
Het-Linear 0.01587 0.01438 0.00723 0.01148 0.05123 0.01391
Het-MLP 0.04397 0.04000 0.03974 0.05053 0.06654 0.04581
Het-SVR 0.03433 0.02250 0.05526 0.03135 0.06141 0.07091
Het-RF 0.02107 0.01608 0.01014 0.01645 0.05892 0.02255
Het-LGBM 0.01934 0.01517 0.01056 0.01407 0.05555 0.01890
Het-XGBoost 0.02116 0.01635 0.01016 0.01595 0.05970 0.02229
Het-CatBoost 0.01748 0.01432 0.00888 0.01262 0.05232 0.01839
@@
Het-Linear 0.01850 0.01438 0.00696 0.01887 0.03357 0.01391
Het-MLP 0.04886 0.03994 0.03924 0.05166 0.06719 0.04581
Het-SVR 0.07261 0.03029 0.03749 0.07211 0.06634 0.07091
Het-RF 0.03444 0.01641 0.00973 0.03068 0.05216 0.02255
Het-LGBM 0.02929 0.01526 0.00972 0.02610 0.04359 0.02255
Het-XGBoost 0.03395 0.01656 0.01000 0.02981 0.05219 0.01890
Het-CatBoost 0.02655 0.01443 0.00879 0.02426 0.04645 0.01839
Appendix H Evaluation cost
We list the table to compare the average evaluation time of experimental (simulation-based) evaluation and our computation evaluation as following table 13. We use the open-source data in table 13 to build the benchmark to test the computational evaluation methods.
@@ Appendix H Evaluation cost
Table 13: Expected evaluation time comparison in 12 scenes per agent.“S” means simulation.
Appendix I Upper bounds tables
@@
Appendix J Paradigm of computational evaluation
@@
Figure 7: The decision tree that how to use computational evaluation models in your own evaluation task. The GEE means generalized evaluation error and the GCEE means generalized causal evaluation error. C is evaluation condition, S is evaluation subject, and M is evaluation metric. $P(M(S)|EC)$ is the potential outcome of evaluation metric after deploying the subject given evaluation condition.